OpenStudy (april):

im doing rules for exponent and radical and its asking me to solve this but i don't understand 15x^3y^-8/5x^2y^-4

OpenStudy (baru):

\[\frac{15 x^3 y^{-8}}{5 x^2 y^{-4}}\]

OpenStudy (baru):

the question is asking you to simplify the above expression as far as possible by using the rules for exponents

OpenStudy (baru):

are you familiar with this rule? \[\frac{a^m}{a^n}=a^{m-n}\]